# Uncovering Autoregressive LLM Knowledge of Thematic Fit in Event Representation

### v2

@@ Abstract
We show closed models possess much thematic fit knowledge and set a new state of the art, while open models also seem to capture much relevant knowledge (in semantic filtering), but yield lower scores. Surprisingly, multi-step reasoning only helped closed models (with few exceptions); generated sentences hurt closed models’ performance; and output form had little to no effect. We analyze the reasons for these findings, and conclude that more foundational work is needed for a single LLM to perform the best on all tasks with the same experimental condition, let alone improve results further. Source code is available at: https://github.com/SafeyahShemali/LLM_Thematic_Fit_25
@@ 1 Introduction
Do large language models (LLMs) have consistent, expressible knowledge (including linguistic knowledge) of event arguments’ thematic fit to their assigned semantic role for a given predicate? A naive approach of prompting LLMs might yield negative or weak results. However, we hypothesize that this knowledge does exist, at least to a certain degree, but needs eliciting. We test our hypothesis along three axes: (1) Reasoning: Does structured reasoning in prompting improve the model’s estimations?111Following chain-of-thought (wei2022chain). (2) Input: Does providing generated sentences as contexts help, compared to providing only a triplet of keywords?222Traditionally consisting of the predicate’s lemma, the argument’s syntactic head’s lemma, and its semantic role. (3) Output: Does specifying categorical output (such as “low” or “high”) yield better results than a numerical output as the model’s thematic fit estimation?
@@
The thematic fit estimation task measures the compatibility between the predicate (typically a verb), its argument (typically a noun phrase), and a specific semantic role assigned to the argument.333Thematic fit is related to, but separate from, semantic role labeling (SRL). Given a sentence, SRL should output the semantic frames in the sentence, including the predicates, arguments, and their semantic roles in the frame. Thematic fit takes the predicates, arguments, and roles as input, and outputs a score. Understanding the thematic fit of individual arguments/roles is thought to be required for SRL, but is not usually modeled explicitly in SRL systems. For theoretical and historical details, see gruber1965studies; fillmore1968lexical; dowty1991thematic; parsons:1990; mcrae1998modeling inter alia. For example, the verb ‘eat’ invokes a set of potential arguments for specific semantic roles. If the role is Agent (usually denoted as Arg0 in PropBank; see sec.§3.2), then ‘people’, ‘cat’, etc., would be a good fit, as opposed to ‘pizza’, ‘apple’, etc., as the latter group would better fit a Patient role (often denoted as Arg1 in PropBank). If the role is Location, then ‘restaurant’, ‘kitchen’, etc., would be a good fit due to their properties. If the role is Instrument, then ‘fork’, ‘knife’, etc. would be a good fit due to their use as tools while eating. And so on. Psycholinguistic experiments show that humans rely on generalized, prototypical knowledge about events and their typical participants during language understanding, early in the process and prior to syntactic interpretation (mcrae2009people; bicknell2010effects).
Thematic fit estimation applies data-driven event representations and computational methods to mimic, or align with this type of human knowledge of events. This is challenging due to the lack of direct labeled training data. Thematic fit norms (see §3.2), which summarize human judgments (aka ratings), provide a benchmark for evaluation, but are not large enough for direct supervised training. Instead, thematic fit models are traditionally trained on proxy data (either PropBank (kingsbury2002treebank) or FrameNet (johnson2016framenet) annotations). Recently, pretrained language models have shown state-of-the-art performance without task-specific supervised training on many computational semantics tasks, such as synonym judgment (levy2017semantic), similarity judgments (hill2015simlex), and more. However, on thematic fit, these models, including masked LMs such as BERT (devlin2019bert), have so far only yielded modest improvements over previous approaches (pedinotti2021cat). In this work, we examine if more recent, large autoregressive models, such as GPT (classicalGPT; radford2018improving; achiam2023gpt), can produce better results via prompting. We attempt to elicit their relevant linguistic knowledge along the following three axes:
@@
Several studies show LLMs perform remarkably well in challenging NLP tasks such as Planning (wang2022towards), automatic scoring (lee2024applying), and question answering (wang2023keqing), when using ‘chain-of-thought prompting’ (wei2022chain). This technique breaks a problem into sub-steps toward answering the core question. It seems to help LLMs in a similar way a scratchpad would help someone solve a math exercise.444It proved so influential that several leading labs have developed ‘thinking’ models with ’test-time compute’, incorporating this approach in the post-training and settings. We view thematic fit estimation as a linguistic reasoning task because it can be broken into clear sub-steps (see §3.1). Thus, we compare the LLMs’ performance using a “naive” single-prompt approach (Simple Prompting hereafter) with the chain-of-thought approach (Step-by-Step Prompting hereafter) on the thematic fit estimation task.
@@
The psycholinguistic data comes in a lemma tuple list format containing a predicate lemma, a lemma of the syntactic head of the predicate’s argument, its thematic role, and the argument’s thematic fit score to its semantic role with this predicate (an average of several human raters’ scores). For example, predicate: eat, argument: pizza, role: Arg0/Agent, score:1.3. Since LLMs were trained on full sentences (as opposed to lemma tuples), this format will be suboptimal for them, even if they possess the relevant linguistic and/or world knowledge. Therefore in some experiments we reconstruct full sentences with these predicates, arguments, and roles.555Upon a close reading, we find that the human raters were indeed provided full sentences for the rating (mcrae1998modeling), but this fact seems to have been lost over the years, and only the lemma tuple format has been used (tilk2016event; hong2018learning; muthupari2022s). We close here a full circle, bringing the conditions closer to the original. See Axis 2 in §3.1 for details.
@@
The thematic fit ratings in our test sets are Likert scale ratings (following mcrae1998modeling) averaged over 7 human raters. Thus, a straightforward evaluation would prompt the model to output numeric scores as well. However, LLMs were designed to handle textual data, so outputting numeric scores may lead to inconsistent results. Moreover, the LLMs’ numeric scores are actually a textual output, which we then convert to its corresponding numeric value. Therefore, we also prompt the LLMs to rate the thematic fit according to predefined categories (“Low”, “High”, etc.), which we then convert to corresponding predefined numeric values for evaluation purposes (see Section §3.2).666We leave for future work whether it is useful to use LLMs’ logprobs for this task, and how to best do that.
@@
Comparison of three axes: reasoning form, input form, and output form.
A novel semantic filtering process to exclude semantically incoherent or semantically incompatible generated sentences.
Showing and analyzing the weakness of new vs. older GPT models in this task, and why chain-of-thought is not always helpful here.
@@ 2 Related Work
Early distributional approaches to the thematic fit task compute the similarity between vector representations of an argument (usually for the argument’s syntactic head word, obtained from a distributional semantic model or static word embeddings) and a prototype vector for the predicate/role pair (erk2010flexible; baroni2010distributional; santus2017; sayeed2015exploration; greenberg2015improving). Our first baseline (hereafter BG) is the distributional approach by greenberg2015improving, which is based on the syntactically structured distributional semantic model by baroni2010distributional. More recent approaches use neural networks to learn event representations. tilk2016event describe a non-incremental role filler model (NN RF, baseline B0) trained to predict role fillers, given as input a predicate, target role, and optionally other contextual roles and their fillers. hong2018learning improve over tilk2016event by adding role prediction as a multi-task objective, and improving pooling of context roles/fillers (ResRoFA-MT, baseline B1). Marton2021ThematicFB replicate hong2018learning’s model and train it on higher quality silver labels, experimenting with the quality/quantity tradeoff of annotations (baseline B2). muthupari2022s focus on evaluating the capability of various models, or model elements, to capture thematic fit. They compare initializing the hong2018learning model with random vectors vs. pre-trained GloVe embeddings (pennington2014glove), and other modifications. We add three of their higher performing models to our baselines: (B3a) Glove-shared-not-tuned, (B3b) Glove-shared-tuned, (B3c) Glove-shared-tuned-20%M. B3a and B3b report the average scores over several runs, whereas B3c is identical to B3b except for reporting the maximum score over several runs, all trained on the same 20% subset of the data.
@@
Contemporary work using pretrained LLMs (without any additional supervised training) for thematic fit either examines if these models can differentiate plausible from implausible sentences (’the teacher bought the laptop’ vs. ’the laptop bought the teacher’) (kauf2023event), or used BERT-style models to predict role fillers for a masked argument in the context of a generated sentence metheniti2020relevant; pedinotti2021cat. Notably, pedinotti2021cat only report a moderate improvement over a distributional approach that uses a high quality structured distributional model (chersoni2019structured). Finally, testa2023we examine the interaction between thematic fit and the ability of LLMs to resolve verbal ellipsis.
We are the first to use prompting with autoregressive LLMs directly for thematic fit.
@@
| Simple Prompting | Head Lemma Tuples | Num. Score | 1.1 |
| Cat. Score | 1.2 | ||
| Generated Sentences | Num. Score | 2.1 | |
| Cat. Score | 2.2 | ||
| Step-by-Step Prompting | Head Lemma Tuples | Num. Score | 3.1 |
| Cat. Score | 3.2 | ||
| Generated Sentences | Num. Score | 4.1 | |
| Cat. Score | 4.2 |
@@ 3 Experiments
The experiments were designed to investigate three axes (§1) with two possible values each. Therefore, we conduct experiments. We refer to them as ’Exp ’ (see Table 1). Each experiment has a prompting template according to the reasoning, input, and output settings, as detailed in Appendix §A. Consistent with the order in the introduction, we will next detail the design for the reasoning form, the input form, and the output form.
3.1 Method: Reasoning, Input and Output
@@
The first axis is concerned with the effectiveness of Step-by-Step Prompting (Exp.3.x and 4.x) compared to Simple Prompting (Exp.1.x and 2.x) in evaluating the thematic fit (leftmost column in Table 1). In the Simple Prompting approach, the model is asked directly only one prompt, which we call a ‘simple thematic fit prompt’, to score the thematic fit of some predicate’s argument to its assigned semantic role (where the predicate, the argument, and the semantic role are all given).
Step-by-Step Prompting decomposes the task into multiple steps, each inferring properties needed to be considered for scoring the thematic fit of the given tuple predicate, argument, semantic role. These steps are: R1. listing the argument’s (salient or relevant) properties, R2. listing the (prototypical or necessary) properties required for the argument in the role “assigned” by the given predicate, R3. listing the (suitable or likely) roles for the argument given predicate, R4. listing the argument’s properties that fit well or are missing for the specified role given the predicate, 5. and finally, estimating the argument’s thematic fit to the assigned role, given these four lists. The last step is similar to the ‘simple thematic fit prompt’ in the Simple Prompting setting. The schematic diagram (Fig. 1) demonstrates the difference between the two approaches. The prompts are given in §A.
@@
The input in each experiment can take one of two forms. The basic form is a tuple predicate, argument, semantic role where the predicate is given in its representative lemma form (verb lemma), and the predicate’s argument is represented by the lemma form of its syntactic head. Adjuncts, subordinate clauses, etc., are ignored. The semantic role is taken from either PropBank or VerbNet. For example, from “I ate a pizza with my friends” we can extract eat, I, Agent or eat, pizza, Patient. Providing more context, and specifically context similar enough to what the LLM presumably was trained on, is what this axis is about. The Lemma Tuples setting (Exp.1.x and 3.x) is how the public test sets are provided, and is also what several previous papers used, including our baselines777even though in the original mcrae1998modeling study, the human raters were given full sentences, but these sentences were not made public as far as we know; see §3.2. In the Generated Sentences setting (Exp.2.x and 4.x), a sentence generation step happens before the other prompts. Sentence generation consists of two steps: 1. Generation: prompting the LLM to reconstruct full sentences with the given predicates, arguments, and roles (see §A.2), and 2. Filtering: instructing the LLM to confirm each generated sentence’s semantic cohesion and its use of the specified predicate and argument in its specified semantic role. The second step is needed as the models (as tested in preliminary trials) seem to often generate semantically incoherent sentences, or use the given argument not in the given semantic role. Each sentence is given three trials to pass the filtering process, which consists of 4 waterfall Yes/No steps: (2a) is the given sentence semantically coherent, (2b) does the sentence contain the predicate, (2c) does the sentence contain the argument as the given predicate’s argument, and (2d) is the argument in the required role for the predicate (see §A.2).
The sentence must pass all the steps of the filtering process in at least one trial, or it will be filtered out. For each of the five sentences, the final fit score is averaged. If steps (1) or (2) fail, we back off to using the lemma tuples. The second column in Table 1 specifies the form of the input.
@@ 3.2 Datasets and Evaluation
We used pre-trained LLMs (see sec. §3.4) “off the shelf” without modification. In an ideal world, we would be able to point the reader to the full pre-training data. We did not add psycholinguistic (or other) labeled data to the training set, nor did we fine-tune the LLMs.
@@
We used the following datasets to evaluate our methods.888These datasets are too small to be split to training, validation and test sets, as is common in machine learning. We follow here previous thematic fit estimation literature. Each of these test sets comprises of a list of tuples predicate, argument, semantic role, score. For example, eat, restaurant, Location, 0.9. The first two are the representative lemma form of the syntactic head of the predicate and argument (in some sentence), the semantic role is as listed below, and the score is the ground truth (human rating). These tuples are given to the LLMs as input, except for the score, which is used to evaluate the models’ output (details below).
mcrae1998modeling, hereafter ’McRae’, contains predicate-argument-role-rating tuples, developed with the participation of 120 native English speakers undergraduate students at the University of Rochester. Students were asked to rate the role typicality (thematic fit) of a predicate (verb) and the argument (noun) in sentence fragments on a Likert scale of 1 to 7. The public dataset does not contain the sentence fragments. The roles are PropBank’s Arg0 and Arg1 (kingsbury2002treebank). We removed eight duplicate records.
pado2006modelling, hereafter ’Pado’, contains 414 tuples in the same form as mcrae1998modeling; however, the role could also be Arg2. To prevent bias in rating the same verb-noun with different roles, one hundred native English speaker volunteers judged tuples in four randomly ordered lists.
ferretti2001integrating evaluate thematic fit for the roles Instrument (hereafter, the Fer-Ins dataset) and Location (hereafter, Fer-Loc dataset). These datasets were done with the input from fifty-eight students. The rating scale is the same as in mcrae1998modeling. However, for Instrument, the participants were asked, “How common is it for someone to use each of the following to perform the eating action?” For Location, the question was, “How common is it for someone to eat in each of the following locations?”. We removed the indefinite articles attached to the argument in the Fer-Loc dataset, in order to consider only the lemma of the syntactic head of the argument (not the phrase).
@@
For each test set, we used Spearman’s Rank Correlation Coefficient () (spearman:1904) between two lists: the test set, sorted by the average human ratings for each item, and the corresponding model output (sorted as well). Categorical output was converted to numeric values, so it can be sorted similarly too before conducting this evaluation (see §3.1).
3.3 Hyperparametes

### v3

@@ Abstract
The thematic fit estimation task measures semantic arguments’ compatibility with a specific semantic role for a specific predicate. We investigate if LLMs have consistent, expressible knowledge of event arguments’ thematic fit by experimenting with various prompt designs, manipulating input context, reasoning, and output forms. We set a new state-of-the-art on thematic fit benchmarks, but show that closed and open weight LLMs respond differently to our prompting strategies: Closed models achieve better scores overall and benefit from multi-step reasoning, but they perform worse at filtering out generated sentences incompatible with the specified predicate, role, and argument.
@@ 1 Introduction
Do large language models (LLMs) have consistent, expressible knowledge of event arguments’ thematic fit to a semantic role for a given predicate?
@@
is the level of compatibility between the predicate (typically a verb), its argument (typically a noun phrase), and a semantic role assigned to the argument. For example, the verb ‘eat’ invokes a set of potential arguments for specific semantic roles. If the role is Agent (or Arg0 in PropBank; see Section §3.2), then ‘people’, ‘cat’, etc., would be a good fit, as opposed to ‘pizza’, ‘apple’, etc., as the latter group would better fit a Patient role (often denoted as Arg1 in PropBank). If the role is Location, then ‘restaurant’, ‘kitchen’, etc., would be a good fit due to their properties. If the role is Instrument, then ‘fork’, ‘knife’, etc. would be a good fit due to their use as tools while eating. Psycholinguistic experiments show that humans rely on such generalized, prototypical knowledge about events and their typical participants during language understanding, early in the process and prior to syntactic interpretation (McRae and Matsuki, 2009; Bicknell et al., 2010).
Thematic fit estimation (TFE)
applies computational methods and event representations derived from data to mimic this type of human knowledge. This is challenging due to the lack of direct labeled training data. Thematic fit norms (see §3.2) summarize human ratings and provide a benchmark for evaluation, but are not large enough for direct supervised training. Instead, TFE models are traditionally trained on SRL datasets as proxy data.111TFE is related to semantic role labeling (SRL). SRL outputs the semantic frames in the input sentence, using frameworks such as PropBank (Kingsbury and Palmer, 2002) or FrameNet (Johnson et al., 2016). The output includes the predicates, arguments, and their semantic roles in the frame. TFE takes the predicates, arguments, and roles as input, and outputs a score. Understanding the thematic fit of individual arguments/roles is thought to be required for SRL, but is not usually modeled explicitly in SRL systems. For theoretical and historical details, see Gruber (1965); Fillmore (1968); Dowty (1991); Parsons (1990); McRae et al. (1998) inter alia. Recently, pretrained language models have shown state-of-the-art performance without task-specific supervised training on many computational semantics tasks, such as synonym judgment (Levy et al., 2017), similarity judgments (Hill et al., 2015), and more. However, on TFE, these models, including masked LMs such as BERT (Devlin et al., 2019), have so far only yielded modest improvements over previous approaches (Pedinotti et al., 2021). In this work, we examine if autoregressive LLMs, such as GPT (Radford et al., 2018; Achiam et al., 2023), can produce better results via prompting. We design our experiments to elicit linguistic knowledge along the following three axes (see §3.1 for details):
@@
LLMs perform remarkably well in challenging NLP tasks such as Planning (Wang et al., 2022), automatic scoring (Lee et al., 2024), and question answering (Wang et al., 2023), when using ‘chain-of-thought prompting’ (Wei et al., 2022). This technique breaks a problem into sub-steps toward answering the core question. It seems to help LLMs in a similar way a scratchpad would help someone solve a math exercise.222It proved so influential that several leading labs have developed ‘thinking’ models with ’test-time compute’, incorporating this approach in the post-training and settings. We view TFE as a linguistic reasoning task because it can be broken into clear sub-steps. Thus, we compare the LLMs’ performance using a “naive” single-prompt approach (simply asking the model to directly score a given lemma tuple; Simple Prompting hereafter), with the chain-of-thought approach (Step-by-Step Prompting hereafter).
@@
The TFE norm data comes in a lemma tuple format containing a predicate lemma, lemma of the syntactic head of the predicate’s argument, its thematic role, and the argument’s thematic fit score (an average of several human raters’ scores). For example, predicate: eat, argument: pizza, role: Arg0/Agent, score:1.3. Since LLMs were pretrained on text rather than such tuples, we hypothesize that they could benefit from seeing full sentences that use the predicate and argument in the given role. We prompt the LLM to generate such sentences and use a novel approach for filtering out semantically incoherent and incompatible sentences.333Upon a close reading, we find that the human raters were indeed provided full sentences for the rating (McRae et al., 1998), but this fact seems to have been lost over the years, and only the lemma tuple format has been used (Tilk et al., 2016; Hong et al., 2018; Muthupari et al., 2022). We close here a full circle, bringing the conditions closer to the original.
@@
The thematic fit ratings in the TFE norms are Likert scale ratings (following McRae et al. (1998)) averaged over 7 human raters. A straightforward evaluation would prompt the model to output numeric scores as well. However, LLMs were designed to handle textual data, so outputting numeric scores may lead to inconsistent results. Therefore, we also prompt the LLMs to rate the thematic fit according to predefined categories (“Low”, “High”, etc.), which we then convert to corresponding predefined numeric values for evaluation purposes (see Section §3.2).444We leave for future work whether it is useful to use LLMs’ logprobs for this task, and how to best do that.
Our experiments compare the performance of two closed-weight LLMs (GPT4.1 and GPT4-Turbo) and two open LLMs (Llama3.2 and Qwen2.5). We show closed LLMs achieve state-of-the-art results and demonstrate the benefits of Step-by-Step Prompting on the TFE task. In contrast, open LLMs demonstrate worse zero-shot performance on this task, but may have an edge on filtering out generated sentences if incompatible with the specified predicate, role, and argument.
@@
Comparison of prompting techniques along three axes: reasoning form, input form (with/out LLM-filtered generated sentences), and output form.
An analysis of related strengths and weaknesses of various open and closed weight LLMs.
Source code is available at: https://github.com/SafeyahShemali/LLM_Thematic_Fit_25
@@ 2 Related Work
TFE is closely related to Selectional Association Resnik (1996), which is an information-theoretic measure of the difference in the probability that some noun class will appear as a specific syntactic argument of the given predicate, compared to the general probability of this class.555It requires a taxonomy (WordNet) for the noun classes and a syntactically parsed corpus for estimating probabilities. Thematic fit focuses on the cognitive load of understanding an utterance with the given predicate and argument(s) in specific thematic roles, post-hoc. FTE also uses a corpus, but with the advent of LLMs, no explicit parses or noun classes are needed, hence it is more applicable and fine-grained.
@@
Early distributional approaches to the TFE task compute the similarity between vector representations of an argument (usually for the argument’s syntactic head word, obtained from a distributional semantic model or static word embeddings) and a prototype vector for the predicate/role pair (Padó et al., 2007; Erk et al., 2010; Baroni and Lenci, 2010; Santus et al., 2017; Sayeed et al., 2015; Greenberg et al., 2015). Our first baseline (hereafter BG) is the distributional approach by Greenberg et al., which is based on the syntactically structured distributional semantic model by Baroni and Lenci. More recent approaches use neural networks to learn event representations. Tilk et al. (2016) describe a non-incremental role filler model (NN RF, baseline B0) trained to predict role fillers, given as input a predicate, target role, and optionally other contextual roles and their fillers. Hong et al. (2018) improve over Tilk et al. (2016) by adding role prediction as a multi-task objective, and improving pooling of context roles/fillers (ResRoFA-MT, baseline B1). Marton and Sayeed (2021) replicate Hong et al. (2018)’s model and train it on higher quality silver labels, experimenting with the quality/quantity tradeoff of annotations (baseline B2). Muthupari et al. (2022) focus on evaluating the capability of various models, or model elements, to capture thematic fit. They compare initializing the Hong et al. (2018) model with random vectors vs. pre-trained GloVe embeddings (Pennington et al., 2014), and other modifications. We add three of their higher performing models to our baselines: (B3a) Glove-shared-not-tuned, (B3b) Glove-shared-tuned, (B3c) Glove-shared-tuned-20%M. B3a and B3b report the average scores over several runs, whereas B3c is identical to B3b except for reporting the maximum score over several runs, all trained on the same 20% subset of the data.
@@
Contemporary work using pretrained LLMs (without any additional supervised training) for TFE either examines if these models can differentiate plausible from implausible sentences (’the teacher bought the laptop’ vs. ’the laptop bought the teacher’; Kauf et al. (2023)), or used BERT-style models to predict role fillers for a masked argument in the context of a generated sentence Metheniti et al. (2020); Pedinotti et al. (2021). Notably, Pedinotti et al. only report a moderate improvement over a distributional approach that uses a high quality structured distributional model (Chersoni et al., 2019). Finally, Testa et al. (2023) examine the interaction between thematic fit and the ability of LLMs to resolve verbal ellipsis. We are the first to use prompting with autoregressive LLMs directly for thematic fit.
@@
| Simple Prompting | Head Lemma Tuples | Numerical | 1.1 |
| Categorical | 1.2 | ||
| Generated Sentences | Numerical | 2.1 | |
| Categorical | 2.2 | ||
| Step-by-Step Prompting | Head Lemma Tuples | Numerical | 3.1 |
| Categorical | 3.2 | ||
| Generated Sentences | Numerical | 4.1 | |
| Categorical | 4.2 |
@@ 3 Experiments
The experiments were structured along three axes (see §1) with two possible values each. Therefore, we conduct experiments. We refer to them as ’Exp ’ (see Table 1). Figure 1 shows a schematic overview of the different methods. Each experiment has a prompting template according to the reasoning, input, and output settings, as detailed in Appendix §A.
3.1 Method: Reasoning, Input, and Output
@@
The first axis is concerned with the effectiveness of Step-by-Step Prompting (Exp.3.x and 4.x) compared to Simple Prompting (Exp.1.x and 2.x) in evaluating the thematic fit. In the Simple Prompting approach we use a single zero-shot prompt to provide the model with a predicate, argument, and semantic role, and ask it to rate the thematic fit. Step-by-Step Prompting decomposes the task into multiple steps, each inferring properties needed to be considered for scoring the thematic fit of the given tuple predicate, argument, semantic role. These steps are: R1. listing the argument’s (salient or relevant) properties, R2. listing the (prototypical or necessary) properties required for the argument in the role “assigned” by the given predicate, R3. listing the (suitable or likely) roles for the argument given predicate, R4. listing the argument’s properties that fit well or are missing for the specified role given the predicate, R5. and finally, estimating the argument’s thematic fit to the assigned role, given these four lists. The last step is similar to the ‘simple thematic fit prompt’ in the Simple Prompting setting.
@@
The input in each experiment can take one of two forms. The basic form (available in the TFE norm data and used in previous work including our baselines) is a predicate, argument, semantic role tuple, where the predicate is a verb lemma and the argument is represented by the lemma form of its syntactic head. Adjuncts, subordinate clauses, etc., are ignored. The semantic role is taken from either PropBank or VerbNet. For example, for the sentence “I ate a pizza with my friends” we would see eat, I, Agent or eat, pizza, Arg1. The Lemma Tuples setting (Exp.1.x and 3.x) uses only the lemmas, as in previous work666In McRae et al. (1998), the raters were given full sentences, but they were not made public as far as we know.. In the Generated Sentences setting (Exp.2.x and 4.x), a sentence generation step happens before the other prompts. We hypothesise that providing more context, similar to what the LLM presumably was trained on, will improve performance. Sentence generation consists of two steps: 1. Generation: prompting the LLM to generate candidate sentences with the given predicate, argument, and role (see §A), and 2. Filtering: instructing the LLM to confirm each generated sentence’s semantic cohesion and its use of the specified predicate and argument in its specified semantic role. The second step is needed, as the models (as tested in preliminary trials) seem to often generate semantically incoherent sentences, or use the given argument not in the given semantic role. Each sentence is given three trials to pass the filtering process, which consists of 4 waterfall Yes/No steps: (2a) is the given sentence semantically coherent, (2b) does the sentence contain the predicate, (2c) does the sentence contain the argument as the given predicate’s argument, and (2d) is the argument in the required role for the predicate (see §A). The sentence must pass all steps of the filtering process in at least one trial, or it will be filtered out. The final fit score is an average over the remaining sentences’ scores. If steps (1) or (2) fail, we back off to using the lemma tuples.
@@ 3.2 Datasets and Evaluation
We used pre-trained LLMs (see Section §3.4) “off the shelf” without modification. In an ideal world, we would be able to point the reader to the full pre-training data. We did not add psycholinguistic (or other) labeled data to the training set, nor did we fine-tune the LLMs.
@@
We used the following datasets to evaluate our methods.777These datasets are too small to be split to training, validation and test sets, as is common in machine learning. We follow here previous TFE literature. Each of these test sets comprises a list of predicate, argument, semantic role, score tuples. For example, eat, restaurant, Location, 0.9. The first two are the representative lemma form of the syntactic head of the predicate and argument (in some sentence), the semantic role is as listed below, and the score is the ground truth (human rating). These tuples are given to the LLMs as input, except for the score, which is used to evaluate the models’ output (details below).
McRae et al. (1998), hereafter ’McRae’, contains predicate-argument-role-rating tuples, developed with the participation of 120 native English speakers undergraduate students at the University of Rochester. Students were asked to rate the role typicality (thematic fit) of a predicate (verb) and the argument (noun) in sentence fragments on a Likert scale of 1 to 7. The public dataset does not contain the sentence fragments. The roles are PropBank’s Arg0 and Arg1, which are roughly equivalent to Agent and Patient, respectively (Kingsbury and Palmer, 2002). We removed eight duplicates.
Padó et al. (2006), hereafter ’Pado’, contains 414 tuples in the same form as McRae et al. (1998); however, the role could also be Arg2. To prevent bias in rating the same verb-noun with different roles, one hundred native English speaker volunteers judged tuples in four randomly ordered lists.
Ferretti et al. (2001) evaluate thematic fit for the roles Instrument (hereafter, the Fer-Ins dataset) and Location (hereafter, Fer-Loc dataset). These datasets were done with the input from fifty-eight students. The rating scale is the same as in McRae et al. (1998). However, for Instrument, the participants were asked, “How common is it for someone to use each of the following to perform the eating action?” For Location, the question was, “How common is it for someone to eat in each of the following locations?”. We removed the indefinite articles attached to the argument in the Fer-Loc dataset, in order to consider only the lemma of the syntactic head of the argument (not the phrase).
@@
For each test set, we used Spearman’s Rank Correlation Coefficient () (Spearman, 1904) between two lists: the test set, sorted by the average human ratings for each item, and the corresponding model output (sorted as well). Categorical output was converted to numeric values, so it can be sorted similarly too before conducting this evaluation (see §3.1).
3.3 Hyperparameters